If the two voltage measurements are V1= 150+_2%And V2= 100+_ 4% respectively. Then the maximum error in the sum of two voltage measurements is 1)+_2.4 % 2) +_2.6% 3)+_2.8%4)+_3.4%

Grade:

1 Answers

Khimraj
5 years ago
error in V1 =150*2/100 = 3V
error in V2 = 100*4/100 = 4V
error in V1+V2 = 7*100/250 = 2.8%
Hope it clears.